OpenStudy (anonymous):

I have a question for my unit 2 geometry portfolio. My mom and I think that answers A and C are the same. Here's the question: Use the Law of Syllogism to draw a conclusion. If a number is a multiple of 12, then it is a multiple of 6. If a number is a multiple of 6, then it is a multiple of 3. Answers: A. If a number is a multiple of 12, then it is a multiple of 6. B. If a number is a multiple of 3, then it is a multiple of 6. C. If a number is a multiple of 3, then it is a multiple of 12. D. If a number is a multiple of 6, then it is a multiple of 12.

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

A. 12,24,36,... B.3,6,9,... C.3,6,9,... D.6,12,18,.. can you conclude ,what the solution may be.

OpenStudy (anonymous):

My mom and I think that A and c are the same answers, so I think that it is one of them. Im sorry I am horrible at geometry.

OpenStudy (anonymous):

Either A or C.

OpenStudy (anonymous):

I have a few more questions that I am stuck on: 2. Write a proof for the following: Given: BC=5; AB=3x Prove AB=6  3. Write a proof for the following: Given: 2/3 (x-6)=8 Prove: X=18 4. Draw a Venn diagram for the following conditional: If you are in Mrs. Lee's Geometry class then you are a student at MICA.d

OpenStudy (anonymous):

for c are 3,6,or 9 multiples of 12 ?

OpenStudy (anonymous):

2. 2x+1=5 2x=5-1=4 x=4/2=2 AB=3 x=3*2=?

OpenStudy (anonymous):

3. \[\frac{ 2 }{ 3 }\left( x-6 \right)=8\] multiply both sides by 3 \[3 \times \frac{ 2 }{ 3 }\left( x-6 \right)=3\times8\] 2(x-6)=24 divide by 2 \[\frac{ 2\left( x-6 \right) }{ 2 }=\frac{ 24 }{ 2 }\] x-6=12 x=12+6=?